OpenStudy (anonymous):

The square of 9 more than a number is equal to 9 more than the square of the number. What is the number?

OpenStudy (anonymous):

n+9^2=n^2+9? @sid0830

OpenStudy (anonymous):

No bro, it will be \[(x+9)^2=x^2+9\] Did you get it?

OpenStudy (anonymous):

Omg, that's the same thing I wrote. Just different variable

OpenStudy (anonymous):

No..I have whole square on LHS..check out yours

OpenStudy (anonymous):

Oh, I see. Okay, now help me work it out. What's the first step.

OpenStudy (anonymous):

kemist's equation is correct

OpenStudy (anonymous):

i dont see square of 'sum'

OpenStudy (anonymous):

Let the number be x 9 more than the number=x+9 Square of 9 more than the number=(x+9)^2 Square of the number=x^2 9 more than square of then number=x^2+9 Put it all together, we get \[(x+9)^2=x^2+9\]

OpenStudy (anonymous):

We just got that part 10 mins ago. I need to know the first step in how to solve it.

OpenStudy (anonymous):

(x+9)^2=x^2+9 x^2+81+18x=x^2+9 18x=9-81=-72 x=-72/18=-4 Is it clear now?

OpenStudy (anonymous):

x=-4? I got the x^2+81 part but the 18x and such I don't understand.

OpenStudy (anonymous):

(a+b)^2 expansion formula